# Causal Data Fusion for Panel Data without Pre-Intervention Period

### v1

@@ Abstract
Traditional panel data causal inference frameworks, such as difference-in-differences and synthetic control methods, rely on pre-intervention data to estimate counterfactuals, which may not be available in real-world settings when interventions are implemented in response to sudden events. In this paper, we introduce two data fusion methods for causal inference from panel data in scenarios where pre-intervention data is unavailable. These methods leverage auxiliary reference domains with related panel data to estimate causal effects in the target domain, overcoming the limitations imposed by the absence of pre-intervention data. We show the efficacy of these methods by obtaining converging bounds on the absolute bias as well as through simulations, showing their robustness in a variety of panel data settings. Our findings provide a framework for applying causal inference in urgent and data-constrained environments, such as public health crises or epidemiological shocks.
Keywords: Causal Inference; Data Fusion; Panel Data; Equi-Confouding Bias; Synthetic Control
@@ 1 Introduction
The difference-in-difference causal inference framework (Card,, 1990; Angrist and Pischke,, 2008) and its significant extension, the synthetic control framework (Abadie and Gardeazabal,, 2003; Abadie et al.,, 2010, 2015), are well-known panel data methods commonly used in such settings. These methods require access to pre-intervention data from the target unit to match it with units (or combinations of units) from the control group. However, in some applications, such pre-intervention data may not be available. For example, in the case of a sudden shock or emergency, immediate action may be taken by the agencies, and hence, no data will be available from the unit during the emergency before the action was taken. For instance, during the COVID-19 pandemic, interventions were implemented almost instantaneously to curb the spread of the virus. As an example, when COVID-19 vaccines first became available in Massachusetts, vaccine doses were accessible only to English-fluent people who were able to use online appointment scheduling sites and who could drive to mass vaccination centers. For many low-income residents who lacked access to information in their preferred language, to a computer or to a car, COVID-19 vaccines remained out of reach, and efforts of a community-based organization in one city (Chelsea) became exemplary for improving vaccine access for their community members. As the organization was instituted (which is the intervention in the system) from almost the first day, these unprecedented circumstances left researchers with no pre-intervention data from the target units to analyze the impact of the interventions. This lack of pre-intervention data hampers the ability to draw robust conclusions about the effectiveness of the interventions and inform future policy decisions.
@@
To address this challenge, in this work, we propose two data fusion-based causal inference methods for panel data with no pre-intervention period. We consider settings where besides the target domain in which we aim to assess the causal effect, we have access to panel data from the units in an auxiliary reference domain. The reference domain can be on time periods before, concurrent or after the time periods in the target domain. Moreover, the observed covariates in the reference domain can be different from those in the target domain. Our first proposed method, called equi-confounding data fusion, is an extension of the ideas of the difference-in-differences approach, and our second proposed method, called synthetic control data fusion, is an extension of the ideas of the celebrated synthetic control method. As mentioned earlier, both the standard difference-in-differences and synthetic control methods require pre-intervention data. Therefore, our contribution is to formalize sufficient conditions regarding the relation between the data in the reference and target domains that can render the identification and estimation of the causal effect using integration (fusion) of the data from the two domains feasible. These conditions essentially allow us to anchor the counterfactual outcome of the unit of interest in the target domain at that in the reference domain. We show the efficacy of these methods by obtaining bounds on the absolute biases, and provided conditions under which the bounds converge to zero. The use of data fusion for identification and estimation of causal parameters in settings where data from one domain is not sufficient for effective evaluation of causal parameters has gained significant attention in the recent years (Dahabreh et al.,, 2020; Degtiar and Rose,, 2023; Colnet et al.,, 2024). However, to the best of our knowledge, we are the first to study this approach for the challenging setting of panel data with single target unit where latent confounding variables may be present.
@@
The rest of the paper is organized as follows. We provide a formal description of the problem in Section 2 and lay out the basic assumptions of the setting under study. We present our proposed equi-confounding bias data fusion and synthetic control data fusion in Sections 3 and 4, respectively. We evaluate our proposed approaches on synthetic data in Section 5. Our concluding remarks are provided in Section 6.
@@ 3.1 Linear Equi-Confounding
Our first proposed approach in an extension of the ideas of negative controls and the difference-in-differences causal inference frameworks to our data fusion setup. Similar extension appeared in (Ghassami et al.,, 2022) for evaluating the long-term causal effect of treatments via data fusion.
@@
To gain intuition about Assumption 2, suppose the intervention in the target domain does not have any causal effect on the outcome variables in the reference domain. For instance, in the vaccine rollout application, this may be satisfied, if we consider the outcome variable in the reference domain as the vaccination rate for another sub-population, which is not able to equally targeted by the community organization. Another example is the influenza vaccination rate for the same sub-population in the case that the community organization did not campaign for the influenza vaccines. Then, Assumption 2 requires that the discrepancy in the mean of the potential outcomes in the absence of any community organization activities in the target domain mirrors that in the reference domain. For instance, the difference in the mean of flu vaccination rate between the target unit and the average of the rest of the units is the same as the same quantity for COVID-19 vaccination rate if the target unit did not receive any influence from the community organization.
@@ 4 Synthetic Control Data Fusion
We consider weight $w_{i}$ for the $i$-th control unit, $12\leq i\leq J+$ 1, and we define the $J$-dimensional weight vector $w=[w_{2}~{}\cdots~{}w_{J+1}]^{\top}$. Our goal is to choose weights such that the following restrictions are (approximately) satisfied.
@@
The intuition of imposing the above restrictions is as follows. Suppose the above matching requirements are satisfied, and also, the outcomes of the units in the reference and target domains are generated by mechanisms that are mostly only difference in the way they are affected by the observed covariates. In this case, if we are matching on the said observed covariates and the outcomes on a long period of time in the reference domain, then we should have been also matching on any unobserved covariate affecting the outcomes in both domains. For instance, for the application of vaccination rate, people’s overall attitude towards and access to vaccines is an unobserved covariate, which may affect COVID-19 vaccination rate and the influenza vaccination rate similarly. Therefore, if we match on flu vaccination rate for an extended time period, as well as all the observed covariates concerning both COVID-19 vaccination and flu vaccination, then there is statistical confidence that we are matching on the attitude and access variables as well.
@@
We note that in restrictions (4)-(6), we will usually have more equations than unknowns, and hence, these restrictions cannot be exactly satisfied. Therefore, we instead, minimize the miss-match in each restriction. Formally, for $i=1,\dots,J+1$, let $F_{i}:=[F_{i,1}~{}\cdots~{}F_{i,T}]^{\top}$, and $K_{i}:=[F_{i}^{\top}~{}Z_{i}^{\top}~{}X_{i}^{\top}]^{\top}$. Weight vector $w$ can be obtained as
$w=\arg\min_{w}\left\|K_{1}-\sum_{i=2}^{J+1}w_{i}K_{i}\right\|_{2}^{2}.$ (8)
An important concern with the approach above is that if we have a larger $T$ than $d_{t}$, the chosen weight vector $w$ may be mostly influenced by $F_{i}$, resulting in a good matching in the reference domain but poor tracking of the $X_{i}$ in the target domain, and hence, poor matching on the potential outcomes in the target domain. In order to quantify the performances of the matchings, for a given $w$, we define the normalized squared errors (NSEs) of the predictors of $F_{1}$, $Z_{1}$, and $X_{1}$, as
@@
Let $b=[b_{F}~{}b_{Z}~{}b_{X}]^{\top}$ be the vector with positive elements that sum up to one, that determines the budgets that we will spend on minimizing the NSEs for $F_{1}$, $Z_{1}$, and $X_{1}$. This vector can be simply set to $b=[1/3~{}1/3~{}1/3]^{\top}$. Alternatively, we can choose to prioritize minimizing the NSE of $F_{1}$, $Z_{1}$, or $X_{1}$ in case we have knowledge that it is more complex than the other two by increasing its corresponding element in $b$. In general, this can be done in a data driven fashion using estimated NSEs. Given the vector $b$, we choose the weight vector $w$ as follows
$w=\arg\min_{w}b_{F}NSE(F_{1},w)+b_{Z}NSE(Z_{1},w)+b_{X}NSE(X_{1},w).$
The objective function above can be written in a more concise way as follows. Define
$\lambda_{F}(b_{F})=\sqrt{\frac{b_{F}}{T}},\quad\lambda_{Z}(b_{Z})=\sqrt{\frac{b_{Z}}{d_{r}}},\quad\lambda_{X}(b_{X})=\sqrt{\frac{b_{X}}{d_{t}}}.$
We replace $K_{i}$ in (8) by its weighted version $K_{i}^{\lambda}:=[\lambda_{F}(b_{F})F_{i}^{\top}~{}\lambda_{Z}(b_{Z})Z_{i}^{\top}~{}\lambda_{X}(b_{X})X_{i}^{\top}]^{\top}$. Then the objective function can be written as $\|K_{1}^{\lambda}-\sum_{i=2}^{J+1}w_{i}K_{i}^{\lambda}\|^{2}_{2}$.
Finally, similar to the original synthetic control method, we require the weights $w$ to be non-negative and that they sum up to one, that is
@@
The purpose of this restriction is interpretability and avoiding over-fitting. Yet, similar to the standard synthetic control method, we can also prevent over-fitting by regularization (Doudchenko and Imbens,, 2016). With this additional restriction on our weights, our approach for choosing weights is by solving the following optimization problem,
$\begin{array}[]{ll}w^{*}=\arg\min_{w}&\left\|K_{1}^{\lambda}-\sum_{i=2}^{J+1}w_{i}K_{i}^{\lambda}\right\|_{2}^{2},\\ {}\\ \text{~{}~{}~{}~{}~{}~{}~{}~{}~{}s.t.}&\sum_{i=2}^{J+1}w_{i}=1,\\ &w_{i}\geq 0,\quad i=2,\ldots,J+1.\end{array}$ (9)
@@ 5 Simulation Studies
For all $i$, $s$, and $t$, entries of $X_{i}$, $Z_{i}$, and $\mu_{i}$ were generated i.i.d. from Unif$(0,1)$, entries of $\phi_{t}$, $\theta_{t}$, $\varphi_{s}$, and $\vartheta_{s}$ were generated i.i.d. from Unif$(0,10)$. $\rho_{t}$ was generated i.i.d. from Unif$(0,20)$ and sorted in the increasing order. The same process was used to generate $\varrho_{s}$ but with Unif$(0,10)$. $\epsilon_{i,t}$ and $\varepsilon_{i,s}$ were generated i.i.d. from Gaussian distribution $\mathcal{N}(0,2)$ and $\mathcal{N}(0,0.5)$, respectively. Finally, the parameters $\alpha_{1,s}$ corresponding to the causal effect on the target unit were generated i.i.d. from Unif$(2,5)$ and sorted in the increasing order. Fixing the factors in the data generating process (DGP) from the above distributions, in order to evaluate our methods, we generated $M=1000$ datasets from the DGP. In order to see the effect of increasing $T$ from $T=T_{1}$ to a larger value $T=T_{2}$, the first $T_{1}$ factor in the latter are the same as the $T_{1}$ factors in the former. For an estimator $\hat{\psi}_{T}$ operating over $T$ time periods in the reference domain, we estimated the bias using $M$ datasets as $1/M\sum_{m=1}^{M}\hat{\psi}_{T}(D_{m})-\psi_{0}$, where $D_{m}$ is the $m$-th dataset.
For the linear equi-confounding approach, let $F_{0}=\frac{1}{J}\sum_{i=2}^{J+1}{F_{i}}$ and $Y_{0}=\frac{1}{J}\sum_{i=2}^{J+1}Y_{i}$. Note that Assumption 2 does not necessarily hold for the factor model described in Equation 10. For instance, Figure 1(a) demonstrates one realization of the DGP ($T=20$) in which the line connecting $\mathbb{E}[F_{1}]$ to $\mathbb{E}[Y_{1}^{(0)}]$ is not parallel to the line connecting $\mathbb{E}[F_{0}]$ to $\mathbb{E}[Y_{0}]$. Hence, the estimator $\hat{\psi}^{eq1}$ is expected to be biased. However, over extensive simulations (several realizations of the DGP), we observed that with our choice of distributions and parameters, the absolute value of the bias normalized by ground truth casual effect in most of the cases was negligible.
Similar concern holds for the logarithm equi-confounding approach. that is, the Assumption 3 does not necessarily hold, Specifically, the lines connecting $\log(\mathbb{E}[F_{1}])$ to $\log(\mathbb{E}[Y_{1}^{(0)}])$ and $\log(\mathbb{E}[F_{0}])$ to $\log(\mathbb{E}[Y_{0}])$ are not parallel (see Figure 1(b)). However, comparing Figures 1(a) and 1(b), in the corresponding DGP we see smaller violation of the logarithmic equi-confounding assumption.
@@
Moving to the synthetic control method, we first solved the optimization problem in Display 9. To ensure good matching on $F$, our choice for $b$ is $[1/4~{}1/4~{}1/2]^{\top}$. We then used the lsei function from the limSolve package (Karline Soetaert et al.,, 2009) in the software R to solve the optimization problem.
@@
To demonstrate the matching performance from the optimization, we calculated the NSEs for latent confounders, covariates, and outcomes in the reference domain, showing that errors decrease with addition of new information. In Figure 3(a), NSEs of $F$, $Z$, and $X$ decrease as time increases. This indicates that the estimated matches become more accurate with more data points. Note that we do not directly target for a good matching for $\mu$ in the optimization problem. Yet the optimized weights $w^{*}$ yield a good match between $\mu_{1}$ and $\sum_{i=2}^{J+1}w^{*}_{i}\mu_{i}$, as seen in Figure 3(b).
Figure 3(c) provides a comparative visualization of the biases from the synthetic control data fusion and equi-confounding data fusion methods. We present the biases calculated by averaging over 1000 datasets for each value of $T$, along with the interquartile range. Overall, we observed a superior performance from the synthetic control data fusion method. The bias from the synthetic control data fusion method decreases as the value of $T$ increases, eventually approaching zero. This reaffirms the result in Theorem 4 that the bias approaches zero as the number of time periods increases. The logarithmic equi-confounding method also presents a decreasing bias trend. However, there is no decreasing pattern for bias from the linear equi-confounding method.
(a) NSEs of $\mu_{i}$, $Z_{i}$, $X_{i}$, and $F_{i}$.
@@
5.1 Placebo Test

### v2

@@ Abstract
Traditional panel data causal inference frameworks, such as difference-in-differences and synthetic control methods, rely on pre-intervention data to estimate counterfactuals. However, such data may not be available in real-world settings when interventions are implemented in response to sudden events, such as public health crises or epidemiological shocks. In this paper, we introduce two data fusion methods for causal inference from panel data in scenarios where pre-intervention data is unavailable. These methods leverage auxiliary reference domains with related panel data to estimate causal effects in the target domain, overcoming the limitations imposed by the absence of pre-intervention data. We show the efficacy of these methods by obtaining converging bounds on the bias as well as through a simulation study. Our proposed methodology renders causal inference feasible in urgent and data-constrained environments where the assumptions of the existing causal inference frameworks are not met. As an application of the proposed methodology, we study the causal effect of the community organization activity on the COVID-19 vaccination rate among the Hispanic sub-population in the city of Chelsea, Massachusetts.
Keywords: Causal Inference; Data Fusion; Panel Data; Equi-Confounding; Synthetic Control
@@ 1 Introduction
The difference-in-difference causal inference framework (Card,, 1990; Angrist and Pischke,, 2008) and its significant extension, the synthetic control framework (Abadie and Gardeazabal,, 2003; Abadie et al.,, 2010, 2015), are well-known panel data methods commonly used in such settings. These methods require access to pre-intervention data from the target unit to match it with units (or combinations of units) from the control group. However, in some applications, such pre-intervention data may not be available. For example, in the case of a sudden shock or emergency, immediate action may be taken by the agencies, and hence, no data will be available from the unit during the emergency before the action was taken. For instance, during the COVID-19 pandemic, interventions were implemented almost instantaneously to curb the spread of the virus. As an example, when COVID-19 vaccines first became available in Massachusetts, vaccine doses were accessible only to English-fluent people who were able to use online appointment scheduling sites and who could drive to mass vaccination centers. For many low-income residents who lacked access to information in their preferred language, to a computer or to a car, COVID-19 vaccines remained out of reach, and efforts of a community-based organization in one city, the city of Chelsea, became exemplary for improving vaccine access for their community members. As the organization was instituted (which is the intervention in the system) from almost the first day, these unprecedented circumstances left researchers with no pre-intervention data from the target units to analyze the impact of the interventions. This lack of pre-intervention data hampers the ability to draw robust conclusions about the effectiveness of the interventions and inform future policy decisions.
@@
To address this challenge, in this work, we propose two data fusion-based causal inference methods for panel data with no pre-intervention period. We consider settings where besides the target domain in which we aim to assess the causal effect, we have access to panel data from the units in an auxiliary reference domain. The reference domain can be on time periods before, concurrent or after the time periods in the target domain. Moreover, the observed covariates in the reference domain can be different from those in the target domain. Our first proposed method, called equi-confounding data fusion, is an extension of the ideas of the difference-in-differences approach, and our second proposed method, called synthetic control data fusion, is an extension of the ideas of the celebrated synthetic control method. As mentioned earlier, both the standard difference-in-differences and synthetic control methods require pre-intervention data. Therefore, our contribution is to formalize sufficient conditions regarding the relation between the data in the reference and target domains that can render the identification and estimation of the causal effect using integration (fusion) of the data from the two domains feasible. These conditions essentially allow us to anchor the counterfactual outcome of the unit of interest in the target domain at that in the reference domain. A challenge introduced by having more than one domain is that extra caution should be taken in matching the target unit with the control units, as naive extension of the existing matching methods may lead to unsatisfactory match in one or both of the domains. We address this issue by introducing a novel data-driven constrained optimization method for designing matches in our proposed synthetic control data fusion approach. We show the efficacy of these methods by obtaining bounds on the absolute biases, and provided conditions under which the bounds converge to zero. The use of data fusion for identification and estimation of causal parameters in settings where data from one domain is not sufficient for effective evaluation of causal parameters has gained significant attention in the recent years (Dahabreh et al.,, 2020; Degtiar and Rose,, 2023; Colnet et al.,, 2024). However, to the best of our knowledge, we are the first to study this approach for the challenging setting of panel data with single target unit where latent confounding variables may be present.
@@
The rest of the paper is organized as follows. We provide a formal description of the problem in Section 2 and lay out the basic assumptions of the setting under study. We present our proposed equi-confounding data fusion and synthetic control data fusion in Sections 3 and 4, respectively. We evaluate our proposed approaches on synthetic data in Section 5. We apply our proposed methodologies to the application of COVID-19 vaccination rate evaluation in Section 6. Our concluding remarks are provided in Section 7.
@@ 3.1 Linear Equi-Confounding
Our first proposed approach is an extension of the ideas of negative controls and the difference-in-differences causal inference frameworks to our data fusion setup. Similar extension appeared in (Ghassami et al.,, 2022) for evaluating the long-term causal effect of treatments via data fusion.
@@
To gain intuition about Assumption 2, suppose the intervention in the target domain does not have any causal effect on the outcome variables in the reference domain. For instance, in the vaccine rollout application, this may be satisfied, if we consider the outcome variable in the reference domain as the vaccination rate for another sub-population, for example, the Black sub-population, which is not able to be equally targeted by the community organization. Another example is the influenza vaccination rate for the same sub-population in the case that the community organization did not campaign for the influenza vaccines. Then, Assumption 2 requires that the discrepancy in the mean of the potential outcomes in the absence of any community organization activities in the target domain mirrors that in the reference domain. For instance, the difference in the mean of COVID-19 vaccination rate for the Black sub-population between the target unit and the average of the rest of the units is the same as the same quantity for COVID-19 vaccination rate for the Hispanic sub-population if the target unit (i.e., the city of Chelsea) did not receive any influence from the community organization.
@@ 4 Synthetic Control Data Fusion
We consider weight $w_{i}$ for the $i$-th control unit, $12\leq i\leq J+$ 1, and we define the $J$-dimensional weight vector $w=[w_{2}~{}\cdots~{}w_{J+1}]^{\top}$, such that the weights must satisfy
Our goal is to choose weights such that the following restrictions are (approximately) satisfied.
@@
The intuition of imposing the above restrictions is as follows. Suppose the above matching requirements are satisfied, and also, the outcomes of the units in the reference and target domains are generated by mechanisms that are mostly only different in the way they are affected by the observed covariates. In this case, if we are matching on the said observed covariates and the outcomes on a long period of time in the reference domain, then we should have been also matching on any unobserved covariate affecting the outcomes in both domains. For instance, for the application of vaccination rate with the choice of the Hispanic and Black sub-populations as the target and reference domains, respectively, peoples’ overall attitude towards vaccines, socioeconomic status, and access to vaccines are unobserved covariates that may similarly affect COVID-19 vaccination rates across the two sub-populations. Therefore, if we match on vaccination rate from one subpopulation for an extended time period, as well as all the observed covariates concerning for COVID-19 vaccination, then there is statistical confidence that we are matching on the unobserved covariates as well.
@@
We note that in restrictions (4)-(6), we will usually have more equations than unknowns, and hence, these restrictions cannot be exactly satisfied. Therefore, we instead, minimize the miss-match in each restriction. Formally, for $i=1,\dots,J+1$, let $\bar{F}_{i}:=[F_{i,1}~{}\cdots~{}F_{i,T}]^{\top}$, and $K_{i}:=[\bar{F}_{i}^{\top}~{}Z_{i}^{\top}~{}X_{i}^{\top}]^{\top}$. A naive way to obtain weight vector $w$ is as follows:
$\begin{array}[]{ll}w=\arg\min_{w}&\left\|K_{1}-\sum_{i=2}^{J+1}w_{i}K_{i}\right\|_{2}^{2},\\ {}\\ \text{~{}~{}~{}~{}~{}~{}~{}~{}~{}s.t.}&\sum_{i=2}^{J+1}w_{i}=1,\\ &w_{i}\geq 0,\quad i=2,\ldots,J+1.\end{array}$ (8)
This is in fact akin to the methodology used in the original synthetic control method. However, unfortunately, this basic approach does not suffice for our data fusion goal. An important concern with the approach above is that if we have a larger $T$ than $d_{t}$, the chosen weight vector $w$ may be mostly influenced by $\bar{F}_{i}$, resulting in a good matching in the reference domain but poor tracking of the $X_{i}$ in the target domain, and hence, poor matching on the potential outcomes in the target domain. Note that this issue does not occur on the standard synthetic control setting because in that setting, matching on the outcomes help better matching on the observed covariates and vice versa. But in our data fusion setting, matching on observed variables in one domain does not necessarily improve that in the other domain. In order to quantify the performances of the matchings, for a given $w$, we define the normalized squared errors (NSEs) of the predictors of $\bar{F}_{1}$, $Z_{1}$, and $X_{1}$, as
@@
To resolve the issue above, we add an extra constraint to the optimization problem that forces the matching on the observed covariates to be close to the best possible matching. We further incorporate a budgeting vector $b$ into our objective function to achieve more flexibility in the choice of the weights. Formally, let $b=[b_{F}~{}b_{Z}~{}b_{X}]^{\top}$ be a vector with positive elements that sum up to one, that determines the budgets that we will spend on minimizing the NSEs for $\bar{F}_{1}$, $Z_{1}$, and $X_{1}$. This vector can be simply set to $b=[1/3~{}1/3~{}1/3]^{\top}$. Alternatively, we can choose to prioritize minimizing the NSE of $\bar{F}_{1}$, $Z_{1}$, or $X_{1}$ in case we have knowledge that it is more complex than the other two by increasing its corresponding element in $b$. In general, this can be done in a data driven fashion using estimated NSEs. We will next propose an algorithm that chooses the budget vector $b$ in a data-driven approach, and uses it to find an optimal weight vector $w^{*}$.
Our method for choosing the optimal budget vector $b^{*}$ and weights $w^{*}$ is outlined in Algorithm 1. This algorithm takes the observed data as well as two hyper-parameters $\eta_{Z}$ and $\eta_{X}$ as the input. We first calculate baseline weights $w^{*}_{Z}$ and $w^{*}_{X}$ separately by minimizing $NSE(Z_{1},w)$ and $NSE(X_{1},w)$, respectively. These baseline weights serve as reference points for constraining the optimization process, as they lead the best possible NSEs that one can get for $X_{1}$ and $Z_{1}$. Next, we define $B$ as the set of all possible vectors $b$ satisfying $b_{F},b_{Z},b_{X}\geq 0$ and $b_{F}+b_{Z}+b_{X}=1$. For each vector $b\in B$, the algorithm finds $w^{*}(b)$ by solving optimization problem 9. This optimization problem aims to find $w$ which minimizes the weighted sum of NSEs. Importantly, we require the solution to be such that the NSEs for $Z_{1}$ and $X_{1}$ are not far away from the best possible NSEs that one can get for these two variables (using baseline weights $w^{*}_{Z}$ and $w^{*}_{X}$). This ensures that the matching for the covariates will not be compromised, and hence resolves the concern mentioned earlier. Finally, the algorithm picks $b^{*}$ by minimizing $NSE(\bar{F}_{1},w^{*}(b))$, and returns $w^{*}=W^{*}(b^{*})$. Therefore, among all the choices with acceptable match on the covariates, we pick the one with the best match on the outcomes in the reference domain. This choice ensures a good match on unobserved components involved in the generating mechanism of the outcomes.
Input: $(\bar{F}_{i},Z_{i},X_{i})_{i=1}^{J+1}$, and hyper-parameters $\eta_{Z}$, $\eta_{X}$.
Define: $B=\{b=\{b_{F},b_{Z},b_{X}\}:b_{F},b_{Z},b_{X}\geq 0\text{ and }b_{F}+b_{Z}+b_{X}=1\}$.
Calculate:
$w^{*}_{Z}=\arg\min_{w}NSE(Z_{1},w),$ s.t. $\sum_{i}w_{i}=1,\quad 0\leq w_{i}$;
$w^{*}_{X}=\arg\min_{w}NSE(X_{1}.w)$ s.t. $\sum_{i}w_{i}=1,\quad 0\leq w_{i}$.
for $b\in B$ do
       Solve:
$w^{*}(b)$ $=\arg\min_{w}b_{F}\cdot NSE(\bar{F}_{1},w)+b_{Z}\cdot NSE(Z_{1},w)+b_{X}\cdot NSE(X_{1},w)$ (9)
s.t. $\quad\frac{1+NSE(Z_{1},w)}{1+NSE(Z_{1},w^{*}_{Z})}\leq 1+\eta_{Z},\quad\frac{1+NSE(X_{1},w)}{1+NSE(X_{1},w^{*}_{X})}\leq 1+\eta_{X},$
$\quad\sum_{i}w_{i}=1,\quad 0\leq w_{i}.$
end for
Set $b^{*}=\arg\min_{b\in B}NSE(\bar{F}_{1},w^{*}(b))$.
Return $w^{*}=w^{*}(b^{*})$.
Algorithm 1 Synthetic control data fusion algorithm
Remark 1.
The objective function in the optimization problem (9) can be written in a more concise way as follows. Define $\lambda_{F}(b_{F})=\sqrt{b_{F}/T}$, $\lambda_{Z}(b_{Z})=\sqrt{b_{Z}/d_{r}}$, and $\lambda_{X}(b_{X})=\sqrt{b_{X}/d_{t}}$. Using this notation, define $K_{i}^{\lambda}:=[\lambda_{F}(b_{F}^{*})\bar{F}_{i}^{\top}~{}\lambda_{Z}(b_{Z}^{*})Z_{i}^{\top}~{}\lambda_{X}(b_{X}^{*})X_{i}^{\top}]^{\top}$. Then the objective function can be written as $\|K_{1}^{\lambda}-\sum_{i=2}^{J+1}w_{i}K_{i}^{\lambda}\|^{2}_{2}$. This form is closer to the form of the objective function in (8). However, note that as mentioned earlier, we have extra constrains in optimization problem (9) compared to (8).
Remark 2.
@@
In the approach above, in line with the original synthetic control method, we require the weights $w$ to be non-negative and that they sum up to one. The purpose of this restriction is interpretability and avoiding over-fitting. Yet, similar to the standard synthetic control method, we can also prevent over-fitting by regularization (Doudchenko and Imbens,, 2016).
@@ 5 Simulation Studies
For all $i$, $s$, and $t$, entries of $X_{i}$, $Z_{i}$, and $\mu_{i}$ were generated i.i.d. from Unif$(0,1)$, entries of $\phi_{t}$, $\theta_{t}$, $\varphi_{s}$, and $\vartheta_{s}$ were generated i.i.d. from Unif$(0,10)$. $\rho_{t}$ was generated i.i.d. from Unif$(0,20)$ and sorted in the increasing order. The same process was used to generate $\varrho_{s}$ but with Unif$(0,10)$. $\epsilon_{i,t}$ and $\varepsilon_{i,s}$ were generated i.i.d. from Gaussian distribution $\mathcal{N}(0,2)$ and $\mathcal{N}(0,0.5)$, respectively. Finally, the parameters $\alpha_{1,s}$ corresponding to the causal effect on the target unit were generated i.i.d. from Unif$(2,5)$ and sorted in the increasing order. Fixing the factors in the data generating process (DGP) from the above distributions, in order to evaluate our methods, we generated $M=300$ datasets from the DGP. In order to see the effect of increasing $T$ from $T=T_{1}$ to a larger value $T=T_{2}$, the first $T_{1}$ factors in the latter are the same as the $T_{1}$ factors in the former. For an estimator $\hat{\psi}_{T}$ operating over $T$ time periods in the reference domain, we estimated the bias using $M$ datasets as $1/M\sum_{m=1}^{M}\hat{\psi}_{T}(D_{m})-\psi_{0}$, where $D_{m}$ is the $m$-th dataset.
@@
For the linear equi-confounding approach, let $F_{0}=\frac{1}{J}\sum_{i=2}^{J+1}{F_{i}}$ and $Y_{0}=\frac{1}{J}\sum_{i=2}^{J+1}Y_{i}$. Note that Assumption 2 does not necessarily hold for the factor model described in Equation 10. For instance, Figure 1(a) demonstrates one realization of the DGP ($T=20$) in which the line connecting $\mathbb{E}[F_{1}]$ to $\mathbb{E}[Y_{1}^{(0)}]$ is not parallel to the line connecting $\mathbb{E}[F_{0}]$ to $\mathbb{E}[Y_{0}]$. Hence, the estimator $\hat{\psi}^{eq1}$ is expected to be biased. However, over extensive simulations (several realizations of the DGP), we observed that with our choice of distributions and parameters, the bias normalized by ground truth casual effect in most of the cases was negligible. Similar concern holds for the logarithmic equi-confounding approach, that is, Assumption 3 does not necessarily hold. Specifically, the lines connecting $\log(\mathbb{E}[F_{1}])$ to $\log(\mathbb{E}[Y_{1}^{(0)}])$ and $\log(\mathbb{E}[F_{0}])$ to $\log(\mathbb{E}[Y_{0}])$ are not parallel (see Figure 1(b)). However, comparing Figures 1(a) and 1(b), in the corresponding DGP we see smaller violation of the logarithmic equi-confounding assumption.
@@
Moving to the synthetic control method, we notice that the optimization problem in Equation 9 takes the form of a quadratic objective function with quadratic constraints. To solve it, we use the gurobi package (Gurobi Optimization, LLC,, 2024) in the software R. One key point of our methodology is the tradeoff between the matching qualities of the reference and the target domains. This balance is governed by the flexibility in terms of the matching quality we allow by the choice of the parameters $\eta_{Z}$ and $\eta_{X}$, which control the relative importance of matching in the reference and target domains, respectively. A larger $\eta_{Z}$ allows for better matching in the target domain at the expense of the reference domain, whereas a larger $\eta_{X}$ allows for a better alignment in the reference domain but may reduce the quality of the matching in the target domain. This tradeoff can be adjusted by the user based on domain-specific considerations in terms of the association between the covariates and the outcome variables, and estimation priorities. In our implementation, we set both $\eta_{Z}$ and $\eta_{X}$ to 0.1. However, we also performed sensitivity analysis to investigate the sensitivity of our estimated causal effect to the choice of $\eta_{Z}$ and $\eta_{X}$. The result of sensitivity analysis is presented in Figure 5(a). As it can be seen in figure, deviations from the chosen value of $0.1$ either increasing or decreasing $\eta_{Z}$ and $\eta_{X}$-exhibit minimal impact on the causal effects.
@@
To demonstrate the matching performance from our proposed method, we calculated the NSEs for latent confounders, covariates, and outcomes in the reference domain. Importantly, the NSE for $\bar{F}_{1}$ starts from a lower value and then converges to a slightly higher value. This is due to the fact that with shorter time period, we will have over-fitting and the algorithm matches on the errors, but with increased length of the time period, the over-fitting is resolved. Note that, the NSE does not converge to zero, because there will be noise for the outcome variable in every time period and hence one should not expect to have an NSE smaller than the variance of the noise. One the other hand, the NSE decreases for $Z_{1}$ because as the length of the time period increases more matching on $\bar{F}_{1}$ helps better matching on $Z_{1}$. This indicates that the estimated matches become more accurate with more data points. Note that we do not directly target for a good matching for $\mu_{1}$ in the optimization problem. Yet the optimized weights $w^{*}$ yield a good match between $\mu_{1}$ and $\sum_{i=2}^{J+1}w^{*}_{i}\mu_{i}$, as seen in Figure 3(b).
(a) NSEs of $\mu_{i}$, $Z_{i}$, $X_{i}$, and $\bar{F}_{i}$.
@@
Figure 3(c) provides a comparative visualization of the biases from the synthetic control data fusion and equi-confounding data fusion methods. We present the biases calculated by averaging over 300 datasets for each value of $T$, along with the interquartile range. Overall, we observed a superior performance from the synthetic control data fusion method. The bias from the synthetic control data fusion method decreases as the value of $T$ increases, eventually approaching zero. This reaffirms the result in Theorem 4 that the bias approaches zero as the number of time periods increases. The logarithmic equi-confounding method also presents a decreasing bias trend. However, there is no decreasing pattern for bias from the linear equi-confounding method.
@@
(a) Sensitivity analysis from the simulated data.
(b) Sensitivity analysis from the real data implementation.
Figure 5: Sensitivity analysis from simulation and application.
5.1 Placebo Test